OpenStudy (anonymous):

When solving x^2 - 4x = -7 by completing the square what value is added to both sides of the equation?

OpenStudy (anonymous):

coefficient of x divided by 2 and then squared (-4/2)^2=?

OpenStudy (anonymous):

Ok so 4? I work it out and 2 and 4. So how do you know which one it is? @CalebBeavers

OpenStudy (anonymous):

-4/2=-2 -2^2=4

OpenStudy (anonymous):

Ok thanks!

OpenStudy (anonymous):

\[x^2-4x=-7\] \[(x-2)^2=-7+4=-3\] \[x-2=\pm\sqrt{3}i\]\[x=2\pm\sqrt{3}i\]